AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Nov 8 13:54:50 2025
Author
*
:
Subject
*
:
<p>16:34</p> <p>Like Run 73, but particle detector in YR10 now active, and recorded in MBS.</p> <p>Clear effect of target on beam lifetime. Target off -> nearly no counts.</p> <p>Attachment 22 - IPM (no target)</p> <p>Attachment 23 - IPM (target)</p> <p>Attachment 24 - YR10 particle detector (gap in events corresponds to target off period)</p> <p>16.42 DSSSD bias & leakage current - attachments 1-2</p> <p>DSSSD #4 currently conditioning at -20V</p> <p>FEE64 temperatures OK - attachment 3</p> <p>ADC, DISC, PAUSE & RESUME data item stats - attachments 4-7</p> <p>System wide checks OK *except* WR deocder status aida04 - attachment 8</p> <p>WR timestamps OK - attachment 9</p> <p>per FEE64 Rate spectra - attachments 10-12</p> <p>per FEE64 1.8.L spectra - attachments 13-14</p> <p>per FEE64 1.8.W spectra - 20us FSR - attachments 15-16</p> <p>pulser peak width aida06 69 ch FWHM, aida08 67 ch FWHM</p> <p>Merger, TapeServer OK - attachments 17-18</p> <p> </p> <p>ASIC settings 2025Apr08-13.45.30</p> <p> aida01-aida04 slow comparator 0xc</p> <p>aida05-aida12 slow comparator 0x8</p> <p>aida13-aida16 slow comparator 0x64, LEC/MEC 0x1</p> <p>17.23 CRYRING vacuum and temperature - attachments 19-21</p> <p> </p> <p>17.34 routine checks OK</p> <p>target density c. 5e+12, beam 3e+6</p> <p> </p> <p>18.40 changed target x,y position to check correlation between target density and detector counts. New position: x:-1500, y:2400.</p> <p>checks OK.</p> <p>target density c. 3.7e+12, beam 3.1e+6</p> <p>20.00 online analysis - per DSSSD energy spectra - attachment 25</p> <p> t, h and p energy peaks clearly observed DSSSD #1 and #2 - p energy peak less clear in DSSSD #3</p> <p> </p> <p>20.30 checks ok</p> <p>target density increased to c. 5e+12 for some minutes then back at c. 3.7e+12, beam 3.1e+6.</p> <p> </p> <p>21.47 checks ok</p> <p>target density increased to c. 5.7e+12 around 21.35, then back at c. 3.7e+12. beam 3.1e+6.</p> <p> </p> <p>22.30 online analysis - per DSSSD energy spectra - attachment 26</p> <p> </p> <p>22.35 Analysis of data file G22-205/R74_29 - attachment 27</p> <p>no timewarps, all deadtimes << 1%</p> <p> </p> <p>22.45 DSSSD#4 bias -20V -> -40V, leakage current -2.7uA -> -4.09uA</p> <p>23.14 DSSS\D#4 leakage current c. -9uA, return to bias -20V</p> <p> </p> <p>23.21 target density slightly unstable in the last hour, changed nozzle position to x=-2000, y=2500.</p> <p> </p> <p><strong>00:04 </strong></p> <p>Temps OK</p> <p>System wide checks OK</p> <p>Rates OK</p> <p>Merger OK (1400 kb/s)</p> <p>Target oscillating, but looks usable. Need to have another look at temperatures tomorrow morning.</p> <p> </p> <p>01:57</p> <p>Temps OK</p> <p>Rates OK</p> <p>Merger OK</p> <p>Target getting thicker.</p> <p> </p> <p>05:00</p> <p>Temps OK</p> <p>Rates OK</p> <p>Merger OK</p> <p>Target fluctuating</p> <p>7:20 All sysytem checks ok. Target lost density @ 6;55 from 4.6e12 to 2.6e12 - will try moving. Update 7:30 after moving density back up to 4.4e12.</p> <p>8:45 All Checks ok but target is fluctuating - really wants to be around 2e12. Had it up to 5e12 about 25 mins ago. Naughty target tbh.</p> <p>9:25 All checks ok, gone to no storage as now tuning next beam energy.</p> <p>10:30 All system checks OK, still tuning.</p> <p>11:38 All system checks ok, still tuning. Frank entered cave to adjust magnetic field as the beam was not cooling at 90 keV/A</p> <p>12:36 All checks okay, cave still open</p> <p>13:40 All checks okay, tuning beam currently</p> <p>14:50 All Checks OK. Target beam overlap currently happening - scanning range not showing anything. Target back on but not hitting - thickness 1.2e12.</p> <p> </p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.3-7933898